Betsy a retiree, requires $5000 per year in extra income.she has $60000 to invest and can invest in B-rated bonds paying 13% per year or in a certificate of deposit paying 3% per year. How much money should be invested in each to realize exactly $5000 in interest per year?

Answer by Earlsdon(6294)

Let x = the amount Betsy will invest in the B-rated bonds paying 13% per annum.
The remaining amount, $60,000-x, she will invest in the certificate of deposit paying 3% per annum.
We can express the earnings on these two investments, after converting the percentages to their decimal equivalents, as:
0.13x+0.03(60000-x) and this is to earn a total of $5,000
Now we can write the necessary equation to solve for x.
0.13x+0.03(60000-x) = 5000 Simplify.
0.13x+1800-0.03x = 5000 Combine the x-terms.
0.1x+1800 = 5000 Subtract 1800 from both sides.
0.1x = 3200 Divide both sides by 0.1
x = 32000
So Betsy needs to invest $32,000 in the B-rated bonds paying 13% interest and the remainder ($60,000-$32,000 = $28,000) in the certificate of deposit paying 3% interest to obtain $5,000 in interest per annum.